OpenStudy (marie152347xoxo):

Ron buys a lawnmower for $1,500. The salesperson says the value will depreciate about 30% per year over the next few years. However, his neighbor says it is likely to depreciate about $300 per year. Which system could be used to determine when the two depreciation models will give the same value for the lawnmower?

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (dumbcow):

well the 2 models are linear and exponential: Linear model \[V = 1500 -300t\] exponential model \[V = 1500(1-0.3)^t = 1500(0.7)^t\] To find when they give same value, set 2 equations equal and solve for t \[1500(0.7)^t = 1500 - 300t\] Unfortunately you will have to estimate since there is no way to solve this algebraically If this is a calculus course, i would use newtons method for approximating zeros Otherwise graph both functions and estimate where they intersect

OpenStudy (jaedyn):

HI

OpenStudy (bonnieisflash1.0):

need some help